# supplementary

### v1

@@ Abstract
This paper introduces HuLP, a Human-in-the-Loop for Prognosis model designed to enhance the reliability and interpretability of prognostic models in clinical contexts, especially when faced with the complexities of missing covariates and outcomes. HuLP offers an innovative approach that enables human expert intervention, empowering clinicians to interact with and correct models’ predictions, thus fostering collaboration between humans and AI models to produce more accurate prognosis. Additionally, HuLP addresses the challenges of missing data by utilizing neural networks and providing a tailored methodology that effectively handles missing data. Traditional methods often struggle to capture the nuanced variations within patient populations, leading to compromised prognostic predictions. HuLP imputes missing covariates based on imaging features, aligning more closely with clinician workflows and enhancing reliability. We conduct our experiments on two real-world, publicly available medical datasets to demonstrate the superiority of HuLP.
@@
Prognosis Representation learning Human-in-the-loop Interactive Imputation-free
@@ 1 Introduction
While deep learning models are emerging as clinical assistants in prognosis, current approaches face two significant problems in the clinical setting. First, the models leave no space for clinicians to intervene, even when the models are incorrect or less confident, thus limiting the clinicians’ ability to provide valuable inputs or correct the models’ predictions. During inference, the models can benefit from such feedback to improve their overall performance, mimicking how doctors collaborate and refine their assessment based on collective expertise. Presently, there is a gap in established methodologies (e.g., [11, 10, 16]) that enable active human interaction and intervention to refine the model’s predictions of clinical features and improve prognosis.
Second, cancer prognosis is particularly challenging when dealing with incomplete data or covariates, compounded with instances where patient outcomes are censored, i.e., instances for which we do not know the exact event time. Missing covariates may result from incomplete data collection [9], patient non-compliance [12], or technical errors [9], while missing outcomes may arise due to patients discontinuing follow-up visits [16], relocating [17], or withdrawing from a study [17]. Standard practice in AI research relies on naive approaches of imputing missing values, such as statistical measures of central tendency (i.e., mean, median, mode), $k$-nearest neighbor (kNN), or more algorithmic approaches, such as multiple imputation by chained equations (MICE). However, in reality, oncologists facing such cases would rely on radiological imaging to gain more insight into the patients’ conditions [7].
The use of electronic health records (EHR) alone in prognosis often falls short of capturing the complex variability among individuals within and across different medical contexts, especially in static non-temporal EHR datasets. For instance, consider two individuals with identical clinical profiles, both diagnosed with lung cancer; despite sharing similar clinical information, their survival outcomes can exhibit significant disparities. Table 1 highlights several such real-world cases from the ChAImeleon [3] lung cancer dataset. Traditional models trained solely on EHR data struggle to reliably distinguish such variations in survival. Notably, the integration of radiological images – which provides a more decadent manifestation of temporal information including age, smoking status, and tumor texture and characteristics – offers a promising avenue for capturing data dynamics that are often overlooked in static clinical data.
In response to these challenges, we introduce Human-in-the-Loop for Prognosis (HuLP) model, a deep learning architecture designed to enhance the reliability and interpretability of prognostic models in clinical settings. Our main contributions are twofold:
@@
  • Capability of handling both missing covariates and outcomes and extract of meaningful vector representations for prognosis: HuLP is equipped with a robust mechanism for handling missing data, ensuring end-to-end reliability in prognostic predictions. By leveraging patients’ clinical information as intermediate concept labels, our model generates richer representations of clinical features, thereby enhancing prognostic accuracy.
@@
70 Male Ex-smoker cT4 cN3 cM1c 3.50 1
70 Male Ex-smoker cT4 cN3 cM1c 1.20 1
@@
72 Male Ex-smoker X X cM1 15.23 0
72 Male Ex-smoker X X cM1 5.17 0
@@
Figure 1: HuLP is composed of (1) a deep learning encoder that extracts features from medical images; (2) an intervention block that allows human intervention during test time; (3) a classifier that ensures concept alignment of the feature embeddings; and (4) a prognosticator that performs survival prediction. Here, (T1, T2, T3) and (Male, Female) are example clinical concepts obtained from the parent categories T-stage and gender, respectively. Our loss is a combination of the concept loss $L_{1}$ applied on the classifier, and prognosis loss $L_{2}$ applied on the prognosticator.
@@ 2 Methodology
Figure 1 shows the complete architecture of the proposed HuLP model which is comprised of four main components: encoder, intervention block, classifier, and prognosticator.
@@
${loss}_{\text{rank}}=\sum_{i,j}E_{i}\,\mathds{1}\{T_{i}<T_{j}\}\exp\left(\frac{\hat{S}(T_{i}\mid w_{i})-\hat{S}(T_{j}\mid w_{j})}{\sigma}\right)\$ (5)
$N$ is the total number of patients and $e_{i}$ is the index of the event time for observation.
@@ 3.1 Datasets
The prognostic ability of HuLP is assessed by comparing it with conventional benchmarks in analyzing two real-world medical datasets: ChAImeleon [3] and HECKTOR [1, 14]. Below we provide a brief overview of each.
The ChAImeleon [3] lung cancer dataset consists of 320 patient CT scans with EHR. The clinical features include age, gender, smoking status, clinical category (T-stage), regional nodes category (N-stage), and metastasis category (M-stage), with up to 26% missing and 59% censored data. For preprocessing, we combined all missing data labels, i.e. “Unknown”, nan, “cTX”, and “cNX” into the same category “X” for each feature. All cancer sub-stages are combined into their parent stage (e.g. cT1a, cT1b, cT1c are combined into cT1). We use a publicly available segmentation model [6] to restrict the ROI to the lung areas.
HECKTOR2021 [1, 14] is a multi-modal head-and-neck cancer dataset comprising 224 CT and PET scans with EHR. The PET and CT scans are registered to a common origin. The clinical features include center, age, gender, TNM 7/8th edition staging and clinical stage, tobacco and alcohol consumption, performance status, HPV status, and treatment (chemoradiotherapy or radiotherapy only), with up to 90% missing and 75% censored data. Features with over 80% missing data are dropped, and all cancer sub-stages are combined into their parent stage. To standardize the inputs, the scans are preprocessed in the same manner for each dataset via resampling, cropping, and resizing.
@@ 3.2 Implementation Details
We compare HuLP against three deep survival methods: DeepHit [11], Deep-MTLR [4], and Fusion [16]. The EHR experiments are run using mode imputation with two FCs of size 64 each followed by a ReLU activation, batch normalization and dropout with probability 0.1, and a prognosticator using a batch size of 96 and learning rate of $1\times 10^{-2}$. We also compare our method against a variation of DeepHit [11] and DeepMTLR [4] using only imaging data as inputs and DenseNet-121 [8] as the encoder to directly predict survival outcomes. Finally, we combine imaging and EHR data to implement the fusion method in [16], the winner of the HECKTOR [1, 14] challenge, as a multimodal approach. We extract imaging features using DenseNet-121 [8], concatenated with EHR through a late fusion technique. We maintain a constant ratio of patients who experienced each event and those who were censored in each fold. The experiments are repeated with two seeds and five-fold cross-validation.
@@ 4 Results
We report the time-dependent concordance (C-index) of Antollini et al. [2] from the survival curves. Table 2 summarizes our results. EHR without images presents the problem of limited depth and richness of static information. Images without EHR leave the model unguided. HuLP consistently demonstrates statistically significant improvements ($p$-value<0.05) over these methods and remain competitive with Fusion [16]; however the learning of fusion from EHR and image embeddings were disjoint. HuLP distinguishes itself by integrating EHR as an intermediate concept labeling that guides the model towards the relevant features, thus producing rich, disentangled embeddings of the clinical features from the images with two added advantages: it allows human expert intervention during test time and is robust to missing data.
To emulate human intervention, $p$ is fully replaced with ground-truth labels for non-missing data, while $p$ is retained for missing data. We run inference on the validation set with and without test-time intervention. Notably, the integration of user interaction and expert intervention of the clinical concepts significantly enhances the model’s prognostic capabilities (Table 3), yielding an improvement of about 0.11 C-index on ChAImeleon [3].
Table 2: Average concordance indices on two seeds and five-fold cross-validation. The highest scores per dataset are bolded. (*) is shown for statistically-significant experiments ($p$-value $<0.05$) based on the average performance of HuLP and the best performing baseline.
ChAImeleon Lung Cancer [3] HECKTOR [1, 14]
DeepHit [11] 0.6522* $\pm$ 0.0371 0.6054* $\pm$ 0.1047
DeepMTLR [4] 0.6624* $\pm$ 0.0643 0.6085* $\pm$ 0.0985
DeepHit [11] (image only) 0.6328* $\pm$ 0.0559 0.7144 $\pm$ 0.0269
DeepMTLR [4] (image only) 0.6400* $\pm$ 0.0361 0.6222* $\pm$ 0.0788
Fusion [16] 0.7399 $\pm$ 0.0534 0.7012 $\pm$ 0.0457
HuLP (ours) 0.7124 $\pm$ 0.0533 0.7329 $\pm$ 0.0415
Table 3: Effect of test-time concept interventions on concordance index scores for ChAImeleon [3]. Results shown are for five-fold cross-validation with two seeds.
@@ 5 Discussion
To our knowledge, HuLP is the first prognostic model that allows human interaction and intervention of known concepts for prognosis. This innovation represents a significant advancement particularly in prognosis, where predicting long-term outcomes can often be more challenging than diagnosing present conditions. Compared to methods where human experts are passive users, HuLP empowers clinicians to actively engage with the model, refining its predictions based on their domain expertise. This collaborative approach fosters a synergistic relationship between humans and computers, allowing each to leverage their strengths. Clinicians, with their deep understanding of clinical features, can provide refined adjustments to the model’s predictions regarding the presence or absence of certain concepts, while HuLP dynamically incorporates these inputs to enhance the accuracy of prognostic assessments. This active collaboration not only improves the interpretability and reliability of prognostic models but also instills confidence in their use in clinical decision-making.

### v2

@@ Abstract
This paper introduces HuLP, a Human-in-the-Loop for Prognosis model designed to enhance the reliability and interpretability of prognostic models in clinical contexts, especially when faced with the complexities of missing covariates and outcomes. HuLP offers an innovative approach that enables human expert intervention, empowering clinicians to interact with and correct models’ predictions, thus fostering collaboration between humans and AI models to produce more accurate prognosis. Additionally, HuLP addresses the challenges of missing data by utilizing neural networks and providing a tailored methodology that effectively handles missing data. Traditional methods often struggle to capture the nuanced variations within patient populations, leading to compromised prognostic predictions. HuLP imputes missing covariates based on imaging features, aligning more closely with clinician workflows and enhancing reliability. We conduct our experiments on two real-world, publicly available medical datasets to demonstrate the superiority and competitiveness of HuLP. Our code is available at https://github.com/BioMedIA-MBZUAI/HuLP.
@@
Prognosis Survival analysis Interactive Human-in-the-loop
@@ 1 Introduction
While deep learning models are emerging as clinical assistants in prognosis, current approaches face two significant problems in the clinical setting. First, the models leave no space for clinicians to intervene, even when the models are incorrect or less confident, thus limiting the clinicians’ ability to provide valuable inputs or improve the models’ predictions. Related works allowing human intervention [1, 5] are applied to natural images but not used for prognosis. During inference, the models can benefit from such feedback to improve their overall performance, mimicking how doctors collaborate and refine their assessment based on collective expertise. Presently, there is a gap in established methodologies (e.g., [13, 12, 18]) that enable active human interaction and intervention to refine the model’s predictions of clinical features to improve prognosis.
Second, in cancer prognosis, dealing with incomplete data and censored patient outcomes (i.e., instances for which we do not know the exact event time) is challenging. Missing covariates may result from incomplete collection [11], non-compliance [14], or technical errors [11], while missing outcomes may arise due to patients discontinuing follow-up visits [18], relocating [19], or withdrawing from a study [19]. Standard practice in AI research typically uses naive imputation methods such as statistical measures of central tendency (i.e., mean, median, mode), $k$-nearest neighbor, or more algorithmic approaches, such as multiple imputation by chained equations (MICE). However, in reality, oncologists rely on radiological images to gain more insights into the patients’ conditions [9].
The use of electronic health records (EHR) alone in prognosis often falls short of capturing the complex variability among individuals within and across different medical contexts, especially in static non-temporal EHR datasets. For instance, consider two individuals with identical clinical profiles, both diagnosed with lung cancer; despite sharing similar clinical information, their survival outcomes can exhibit significant disparities. Table 1 highlights several such real-world cases from the ChAImeleon [4] lung cancer dataset. Traditional models trained solely on EHR data struggle to reliably distinguish such variations in survival. Notably, the integration of radiological images – which provides a richer manifestation of temporal information, including age, smoking status, and tumor texture and characteristics – offers a promising avenue for capturing data dynamics that are often overlooked in static clinical data.
In response to these challenges, we introduce Human-in-the-Loop for Prognosis (HuLP), a deep learning architecture inspired by [1, 5] designed to enhance the reliability and interpretability of prognostic models in clinical settings. Our main contributions are twofold:
@@
  • Capability of handling both missing covariates and outcomes and extraction of meaningful vector representations for prognosis: HuLP is equipped with a robust mechanism for handling missing data, ensuring end-to-end reliability in prognostic predictions. By leveraging patients’ clinical information as intermediate concept labels, our model generates richer representations of clinical features, thereby enhancing prognostic accuracy.
@@
70 Male Ex-smoker T4 N3 M1c 3.50 1
70 Male Ex-smoker T4 N3 M1c 1.20 1
@@
72 Male Ex-smoker X X M1 15.23 0
72 Male Ex-smoker X X M1 5.17 0
@@
Figure 1: HuLP is composed of (1) a deep learning encoder that extracts features from medical images; (2) an intervention block that allows human intervention during test time; (3) a classifier that ensures concept alignment of the feature embeddings; and (4) a prognosticator that performs survival prediction. Here, (T1, T2, T3) and (Male, Female) are example clinical concepts obtained from the parent categories T-stage and gender, respectively. Our loss is a combination of the concept loss $L_{1}$ applied on the classifier and prognosis loss $L_{2}$ applied on the prognosticator.
@@ 2 Methodology
Figure 1 shows the complete architecture of the proposed HuLP model, which is comprised of four main components: encoder, intervention block, classifier, and prognosticator.
@@
${loss}_{\text{rank}}=\sum_{i,j}E_{i}\,\mathds{1}\{T_{i}<T_{j}\}\exp\left(\frac{\hat{S}(T_{i}\mid w_{i})-\hat{S}(T_{j}\mid w_{j})}{c}\right)\$ (5)
$N$ is the total number of patients, $e_{i}$ is the index of the event time for observation, and $c$ is set to a constant 0.1 following [13].
@@ 3.1 Datasets
The prognostic ability of HuLP is assessed by comparing it with conventional benchmarks in analyzing two real-world medical datasets: ChAImeleon [4] and HECKTOR [2, 16]. Below, we provide a brief overview of each.
The ChAImeleon [4] lung cancer dataset consists of 320 patient CT scans with EHR. The clinical features include age, gender, smoking status, clinical category (T-stage), regional nodes category (N-stage), and metastasis category (M-stage), with up to 26% missing and 59% censored data. For preprocessing, we combined all missing data labels, i.e. “Unknown”, nan, “TX”, and “NX” into the same category “X” for each feature. All cancer sub-stages are combined into their parent stage to increase the number of samples per category (e.g. T1a, T1b, T1c are combined into T1). We use a publicly available segmentation model [8] to restrict the ROI to the lung areas.
HECKTOR [2, 16] is a multi-modal head-and-neck cancer dataset comprising 224 CT and PET scans with EHR. The PET and CT scans are registered to a common origin. The clinical features include center, age, gender, TNM 7/8th edition staging and clinical stage, tobacco and alcohol consumption, performance status, HPV status, and treatment (chemoradiotherapy or radiotherapy only), with up to 90% missing and 75% censored data. Features with over 80% missing data are dropped, and all cancer sub-stages are combined into their parent stage. To standardize the inputs, the scans are preprocessed in the same manner for each dataset via resampling, cropping, and resizing.
@@ 3.2 Implementation Details
We compare HuLP against three deep survival methods: DeepHit [13], Deep-MTLR [6], and Fusion [18]. DeepHit [13] and Deep-MTLR [6] are chosen because they are both top-performing discrete survival methods in prognosis; similarly, our HuLP implementation is also discrete. Fusion [18] is chosen as a multimodal baseline and also because it won the HECKTOR [2, 16] competition, the same dataset used in this work.
DeepHit [13] and Deep-MTLR [6] (EHR) are run using mode imputation with two FCs of size 64 each followed by a ReLU activation, batch normalization and dropout with probability 0.1, and a prognosticator using a batch size of 96 and learning rate of $1\times 10^{-2}$. We also compare our method against a variant of DeepHit [13] and DeepMTLR [6] using imaging data as inputs and DenseNet-121 [10] as the encoder to directly predict survival outcomes. Finally, we implement the idea of Fusion [18] by extracting imaging features using DenseNet-121 [10], concatenated with EHR through a late fusion technique. We maintain a constant ratio of patients who experienced each event and those who were censored in each fold. The experiments are repeated with two seeds and five-fold cross-validation.
@@ 4 Results
We report the time-dependent concordance (C-index) of Antolini et al. [3] from the survival curves. Table 2 summarizes our results. EHR without images presents the problem of limited depth and richness of static information. Images without EHR leave the model unguided. HuLP consistently demonstrates statistically significant improvements ($p$-value<0.05) over these methods and remains competitive with Fusion [18]; however, the learning of fusion from EHR and image embeddings was disjoint. HuLP distinguishes itself by integrating EHR as an intermediate concept labeling that guides the model towards the relevant features, thus producing rich, disentangled embeddings of the clinical features from the images with two added advantages: it allows human expert intervention during test time and is robust to missing data.
To emulate human intervention, $p$ is fully replaced with ground-truth labels for non-missing data, while $p$ is retained for missing data. We run inference on the validation set with and without test-time intervention. Notably, the integration of user interaction and expert intervention of the clinical concepts significantly enhances the model’s prognostic capabilities (Table 3), yielding an improvement of about 0.1 C-index on ChAImeleon [4].
Table 2: Average concordance indices on two seeds and five-fold cross-validation. The highest scores per dataset are bolded. (*) is shown for statistically significant experiments ($p$-value $<0.05$) based on the average performance of HuLP and the best-performing baseline.
Modality ChAImeleon Lung Cancer [4] HECKTOR [2, 16]
DeepHit [13] EHR 0.6522* $\pm$ 0.0371 0.6054* $\pm$ 0.1047
DeepMTLR [6] EHR 0.6624* $\pm$ 0.0643 0.6085* $\pm$ 0.0985
DeepHit [13] Image 0.6328* $\pm$ 0.0559 0.7144 $\pm$ 0.0269
DeepMTLR [6] Image 0.6400* $\pm$ 0.0361 0.6222* $\pm$ 0.0788
Fusion [18] Image+EHR 0.7399 $\pm$ 0.0534 0.7012 $\pm$ 0.0457
HuLP (ours) Image+EHR 0.7124 $\pm$ 0.0533 0.7329 $\pm$ 0.0415
Table 3: Effect of test-time concept interventions on concordance index scores for ChAImeleon [4]. The results shown are for five-fold cross-validation with two seeds.
@@ 5 Discussion
To our knowledge, HuLP is the first prognostic model that allows human interaction and intervention of known concepts for prognosis. This innovation represents a significant advancement particularly in prognosis, where predicting future outcomes can often be more challenging than diagnosing present conditions. Compared to methods where human experts are passive users, HuLP empowers clinicians to actively engage with the model, refining its concept predictions based on their domain expertise. This collaborative approach fosters a synergistic relationship between humans and computers, allowing each to leverage their strengths. Clinicians, with their deep understanding of clinical features, can provide refined adjustments to the model’s predictions regarding the presence or absence of certain concepts, while HuLP dynamically incorporates these inputs to enhance the accuracy of prognostic assessments. This active collaboration not only improves the interpretability and reliability of prognostic models but also instills confidence in their use in clinical decision-making.
@@ 6 Conclusion
{credits}
6.0.1 Acknowledgements
De-identified data used in the development of this solution were prepared and provided by the ChAImeleon Project within the Open Challenges organized by its Consortium, funded by the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 952172, entitled ‘Accelerating the lab to market transition of AI tools for cancer management.